OpenStudy (lgbasallote):

Lgbariddle >:)) Using any Maths Symbols or Signs prove it... 0 0 0 = 6 1 1 1 = 6 2 2 2 = 6 3 3 3 = 6 4 4 4 = 6 5 5 5 = 6 6 6 6 = 6 7 7 7 = 6 8 8 8 = 6 9 9 9 = 6 You can Insert Everything but Numbers, And you can't change everything.

OpenStudy (anonymous):

Can I insert any amount of symbols?

OpenStudy (anonymous):

Lgbariddle- nice name.

OpenStudy (blockcolder):

It's easy for 5: 5+5/5=6, but I have no idea how to make 6 out of 0's. =))

OpenStudy (anonymous):

I will do a few: (1+1+1)! = 6

OpenStudy (lgbasallote):

yes..as long as you dont put new numbers in

OpenStudy (lgbasallote):

1 down :D

OpenStudy (apoorvk):

(0+0+0)!=6 (1+1+1)!=6 2+2+2=6 3*3 -3 = 6 sqrt4 + sqrt4 +sqrt4 = 6 next batch coming up

OpenStudy (anonymous):

That's works for (0!+0!+0!)!=6

OpenStudy (apoorvk):

oops 0 i did wrong

OpenStudy (apoorvk):

LOL o.O

OpenStudy (anonymous):

\[0+0+0+\frac{x+x+x+x+x+x}{x}=6\]lolol

OpenStudy (anonymous):

Hey, nobody accused me of being clever. :D

OpenStudy (lgbasallote):

hmmm 0-4 down

OpenStudy (lgbasallote):

lol @badreferences =)))

OpenStudy (anonymous):

\[ (\cos (0) + \cos (0) + \cos (0) )!= 6 \]

OpenStudy (lgbasallote):

^that's actually the real one for 0 hehe

OpenStudy (apoorvk):

well, 0! method works as well.

OpenStudy (lgbasallote):

guess so haha

OpenStudy (anonymous):

\[\left(\lim_{x\to\infty}\frac{9+9+9}{x}\right)+\frac{x+x+x+x+x+x}{x}\]lololol

OpenStudy (lgbasallote):

lol no x's :PPP

OpenStudy (blockcolder):

6+6-6 =6

OpenStudy (blockcolder):

7-7/7=6

OpenStudy (anonymous):

Okay I will do from the behind, \[ (\sqrt9 \times \sqrt9 ) – \sqrt9 = 6 \]

OpenStudy (lgbasallote):

6 and 7 down

OpenStudy (lgbasallote):

8 left

OpenStudy (lgbasallote):

oh 5 too

OpenStudy (anonymous):

Sum of the cube roots of three 8 is 6.

OpenStudy (apoorvk):

i dont thing using a cube root would be allowed, would it?

OpenStudy (anonymous):

wow lgba

OpenStudy (apoorvk):

oh well, ffm did it, so must be

OpenStudy (lgbasallote):

izallowed @apoorvk so 8 down

OpenStudy (lgbasallote):

5 is easy =))

OpenStudy (anonymous):

\sqrt{9}-\sqrt{9}+\sqrt{9}!

OpenStudy (anonymous):

Like seven.

OpenStudy (lgbasallote):

lol that's 2 :P hahahah

OpenStudy (anonymous):

6+6-6 = 0

OpenStudy (anonymous):

\[\sqrt9-\sqrt9+\sqrt9!=3!=6\]

OpenStudy (lgbasallote):

6+6-6 is 6 not zero lol

OpenStudy (anonymous):

so are we done?

OpenStudy (anonymous):

No, I'm lobbying that my previous answers are the best.

OpenStudy (lgbasallote):

yeah =)) but 5 is 5 + 5/5 not 5/5 + 1 lol

OpenStudy (anonymous):

lol i meant 6

OpenStudy (apoorvk):

sqrt9 * sqrt9 - sqrt9

OpenStudy (anonymous):

In particular \[\left(\lim_{x\to\infty}\frac{9+9+9}{x}\right)+\frac{x+x+x+x+x+x}{x}\]

OpenStudy (lgbasallote):

it is finished

OpenStudy (anonymous):

Oh yes typo :P

OpenStudy (anonymous):

3! +3-3= 6

OpenStudy (lgbasallote):

lol badrefs with the x's hahaha

OpenStudy (apoorvk):

oh, nine's done :/

OpenStudy (apoorvk):

Good one, LGBA

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

wait is my 3 one right?

OpenStudy (lgbasallote):

could be....hehe =))

OpenStudy (anonymous):

yeah!